Need help with problem asap
U. the greater c. the smaller momentum d. the greater speed e the greater acceleration 12. Blocks A and B are moving toward each other A has a mass of 2.0 kg and a velocity of 50 m/s while B has a mass of 4.0 kg and a velocity of -25 m/s. They suffer a completely inelastic collision. The kinetic energy lost during the collision is a. 5600 J b. 1250 J d. 3750 J e. 5000 J 13. A 4-kg ball is dropped vertically from a height and it hits the floor with a speed of 10 m/s. Th ball rebounds vertically and reaches a height of 2 m above the floor. The magnitude ot impuls on the ball during the contact with the floor is: a. 78 N.s b. 65 N.s c. 40 N.s d. 15 N.s e. 8 N.s
Answer #1

Answer:(d) 3750J

From the given data

The initial momentum is

2*50-4*25 =0

Then after collision they stick together then the final velocity is zero

Now the loss in kinetic energy is = (1/2)(2)(50)2+(1/2)(4)(25)2 =3750J

13)

Answer is (b)

First velocity of the ball just before impact is (v1) =10m/s

The velocity just after touch the ground is v2 = √2*g*h2 =√2*9.81m/s2*2m=6.264m/s

Now the impulse is nothing but the change in momentum =mv1+mv2 =4(10)+4(6.264) =40+25.056 =65.056N.s